Dr. Gruner brings his old pocket watch to be repaired. The gears stopped and will not turn. What is the probability that the hour hand is between the 5 and the 8?
Answer #1

Total hours in dial = 12

Between 5-8, number of hours =3

So, probability =3/12

=1/4 or 25%